solve the system of equations by using the substitution method
2x-3y=9, y=x-4

Answer by jim_thompson5910(35256)

2x-3y=9


2x-3(x-4)=9


2x-3(x-4)=9


2x-3x+12=9


-x+12=9


-x=9-12


-x=-3


x=(-3)/(-1)


x=3

-------------------------------------------------------

y=x-4


y=3-4


y=-1


So the solutions are x = 3 and y = -1